OpenStudy (mackenzie2013):

Which of the following ratios are equivalent to 6:8? 36/20 9:12 12 to 16 15/20 24:36 16 to 24

OpenStudy (paki):

6:8 = 3/4... and 9:12 = 3/4...

OpenStudy (anonymous):

3;4 ; 6: 8 ; 9:12 ; 12:16 ; 15:20 ;

OpenStudy (mackenzie2013):

Thank you guys! :)